OpenStudy (anonymous):

PLEASE HELP!!! The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2). Answers: y = 2x + 4 y = negative 1 over 2x − 1 over 2 y = − 1 over 2x − 7 over 2 y = 2x − 8

OpenStudy (anonymous):

equation of a line parallel to AB is y=2x + c. now put y,x as 3,-2 and you get the value of c. that's your answer

OpenStudy (anonymous):

i don;t understand how you got that at all the way you worded your answer is too confusing, where did you get c?

OpenStudy (anonymous):

okay first of all the lines are parallel therefore both of them have the same slope(m) and u r given the point (3,-2) just put the slope m=2 and the point(3,-2) in this equation y-y1=m(x-x1) and solve

OpenStudy (anonymous):

D. is the answer

OpenStudy (anonymous):

Y=mx+b so you can see where he got the "m"

OpenStudy (anonymous):

thank you all! I understand

OpenStudy (anonymous):

Parallel lines have the same slopes and you are given coordinates allowing you to plug and chug into the formula